# Quantum emitters in silicon: Purcell-enhanced lifetime modulation as a probe of local refractive index changes

### v2

@@ Abstract
Quantum emitters embedded in silicon photonic cavities offer a powerful and scalable platform for label-free refractive index sensing at the nanoscale. We propose and theoretically analyze a novel sensing mechanism based on Purcell-enhanced modulation of the emitter’s spontaneous emission lifetime, enabling detection of refractive index changes via time-correlated single-photon counting (TCSPC). In contrast to traditional resonance-shift sensors, our approach exploits the lifetime sensitivity to variations in the local density of optical states (LDOS), providing an intensity-independent, spectrally unresolvable, and CMOS-compatible sensing modality. We derive analytical expressions linking refractive index perturbations to relative lifetime shifts and identify an optimal off-resonance operation regime, where the lifetime response becomes linear and maximally sensitive to small perturbations. Parametric evaluation of the analytical expressions shows that, for moderate-quality photonic cavities ($Q$ = 105–107), this method enables refractive index detection limits as low as 10-9 RIU — competitive with or even outperforming state-of-the-art plasmonic and microresonator sensors, yet requiring significantly simpler instrumentation. Furthermore, we show that long-lived emitters such as T-centers in silicon provide a unique advantage, allowing sub-nanosecond lifetime shifts to be resolved with standard TCSPC systems. Although room-temperature operation of quantum emitters in silicon has yet to be demonstrated, our results lay the theoretical foundation for scalable, room-temperature, quantum-enabled sensing of refractive index changes — eliminating the need for spectral resolution and cryogenic infrastructure.
@@ I Introduction
However, their practical use at room temperature remains constrained by phonon-induced dephasing and non-radiative recombination, which broaden and suppress ZPL emission. To address these limitations, we propose a fundamentally different transduction approach: lifetime-based sensing via modulation of the local density of optical states (LDOS). Unlike conventional schemes based on spectral shifts, this method relies on variations in the emitter’s spontaneous emission rate induced by its electromagnetic environment. Such a detection scheme is inherently more robust to thermal broadening and inhomogeneous line shifts, and remains functional even when the ZPL becomes spectrally unresolved. In other words, even if the emitter rapidly loses coherence (i.e., short $T_{2}$) and exhibits a broad spectral line, the excited-state population may still decay on a much longer timescale ($T_{1}$), yielding a measurable lifetime — the quantity we exploit for sensing. This approach extends beyond T-centers to a broader class of T-like color centers, including group III — carbon complexes such as B–C, Al–C, and In–C [7, 8]. These centers are predicted to exhibit similar spin and optical properties, with higher structural symmetry (e.g., $C_{2v}$), improved radiative efficiencies, and greater thermal stability — making them promising candidates for scalable integration with silicon photonics.
@@
The viability of this lifetime-based sensing scheme is further supported by decades of work on quantum spin systems in silicon. For instance, phosphorus donor nuclear spins in isotopically purified silicon have shown coherence times exceeding 39 minutes at room temperature [2], while 29Si nuclear spin qubits maintain coherence even without isotopic enrichment [3], affirming silicon’s status as an ultra-low-noise host for quantum systems. Analogous temperature studies of color centers such as SiV in diamond provide insights into emission linewidths: homogeneous broadening increases with temperature following a $T^{3}$ dependence, leading to ZPL broadening from $\sim$0.1–0.6 nm at cryogenic temperatures to $\sim$1.3–2 nm at room temperature [9]. A similar trend is anticipated for silicon-based centers, though experimental verification remains pending. Although room-temperature operation of quantum emitters in silicon remains unproven, no conclusive physical limitation has been established, suggesting potential for future advancement through photonic or material engineering.
@@
In light of these considerations, our proposed sensing scheme — based on photonic cavity-enhanced lifetime modulation of optically active color centers [10, 11, 12, 13, 14] — remains viable under room-temperature conditions. By decoupling sensing performance from spectral resolution constraints, it offers a robust and scalable platform for refractive index detection and single-molecule interaction studies. It is inherently compatible with the future discovery and optimization of silicon-based quantum emitters.
@@ II Theoretical framework and results
$\tau=\frac{\tau_{0}}{F_{P}},$ (3)
where $\tau_{0}$ is the intrinsic spontaneous emission lifetime of the emitter in free space or in an unstructured (homogeneous) dielectric. We assume that the optical cavity used to modify the spontaneous emission lifetime of the emitter is optimized to achieve the maximum LDOS and Purcell enhancement. Local changes in the refractive index induce a spectral detuning of the cavity resonance. This detuning reduces the LDOS at the emitter’s location, thereby suppressing the spontaneous emission rate and resulting in an increase in the emitter’s lifetime. This lifetime shift can be detected using time-correlated single-photon counting (TCSPC), forming the basis for a label-free sensing modality. Given this, we estimate the effect of a refractive index change on the spontaneous emission lifetime of a quantum emitter coupled to a photonic cavity. A local change in the refractive index alters the effective index of the cavity mode, $n_{\mathrm{eff}}$, leading to a shift in cavity’s resonance angular frequency $\omega_{\mathrm{cav}}$. In the regime of small perturbations, the shift in the resonance frequency can be expressed using first-order perturbation theory as:
[…]
where $\Delta\varepsilon(\vec{r})=2n(\vec{r})\Delta n(\vec{r})$ is the permittivity perturbation induced by the local refractive index change $\Delta n(\vec{r})$, and $\vec{E}(\vec{r})$ is the cavity mode electric field. By assuming $\Delta\varepsilon=2n\Delta n\approx 2n_{\mathrm{eff}}\Delta n$, we can then simplify Eq. (4) to:
@@
where $\Delta\omega=\omega_{\mathrm{emitter}}-\omega_{\mathrm{cav}}$ is the detuning. If the cavity is on-resonance in the initial state: $\Delta\omega=0$, and $F_{P}=F_{P,\mathrm{max}}$, and lifetime is at its minimum ($\tau=\tau_{0}/F_{P}$). Then, a small change in $F_{P}$ leads to the following relative lifetime change of the emitter:
@@
Figure 1: Simulated lifetime change ($\Delta\tau/\tau$) versus refractive index change ($\Delta n$): (a) linear scale for $\Delta\tau/\tau$ from 0 to 10% and logarithmic scale for $\Delta n$; (b) logarithmic scale for $\Delta\tau/\tau$ from $10^{-1}$ to $10^{5}$ and logarithmic scale for $\Delta n$.
@@
However, at resonance ($\Delta\omega=0$), the Purcell factor is at its maximum, but its derivative with respect to detuning is zero. Therefore, in the immediate vicinity of $\Delta n=0$ (i.e., near the point $n_{\mathrm{eff}}$, see Figures 1a and 2), the dependence of $\Delta\tau/\tau$ on $\Delta n$ becomes non-linear (specifically, quadratic; see Eq. (7)), making it unsuitable for applications that require quantitative detection of small refractive index changes.
Figure 2: Simulated normalized Purcell factor versus refractive index change $\Delta n$: (a) linear scale for $\Delta n$ from $-10^{-4}$ to $+10^{-4}$; (b) log scale for $\Delta n$ from $10^{-9}$ to $10^{-3}$.
@@
We next derive the expression for the detection limit related to these optimal detection points for each $Q$ value, $\Delta n_{\mathrm{opt}}$. The emission lifetime $\tau$ at the shifted point $n_{0}$ can be written as:
@@
=±$=\pm 2\tau_{0}\frac{Q}{n_{\mathrm{eff}}}=\pm\tau_{\mathrm{eff}}\frac{Q}{n_{\mathrm{eff}}},$ (9)
here $\tau_{0}=\tau(n_{0})=0.5\tau(n_{\mathrm{eff}})=0.5\tau_{\mathrm{eff}}$. In the vicinity of $n_{0}$, we can write:
@@
Eq. (13) estimates the minimum detectable $\Delta n_{\text{opt}}$ for a given temporal resolution. According to Eq. (13), the detection limit of the proposed refractive index sensing approach in the off-resonance regime, $\Delta n_{\text{opt}}$, is limited by the $Q$ value of the photonic cavity and the resolution capability $\Delta\tau/\tau_{0}$ of the TCSPC system used. For instance, if a TCSPC system can resolve a 3% change in lifetime ($\Delta\tau/\tau_{0}=0.03$), then the minimum detectable $\Delta n_{\text{opt}}$ would be approximately $3.75\times 10^{-7}$ RIU for $Q=10^{5}$, $3.75\times 10^{-8}$ RIU for $Q=10^{6}$, and $3.75\times 10^{-9}$ RIU for $Q=10^{7}$, assuming $n_{\text{eff}}=2.5$. In addition, we calculated the linear dynamic range (LDR) for each $Q$ value as the span of refractive index changes $\Delta n$ over which the relative lifetime change $\Delta\tau/\tau_{0}$ remains approximately linear (with a deviation less than $\pm$10from linear behavior) around the $n_{0}$ point. These data are summarized in Table 2.
@@
$10^{5}$ $3.75\times 10^{-7}$ $1.11\times 10^{-5}$
$10^{6}$ $3.75\times 10^{-8}$ $1.11\times 10^{-6}$
$10^{7}$ $3.75\times 10^{-9}$ $1.11\times 10^{-7}$
@@
where $\left\langle\cos^{2}\theta\right\rangle=\int_{0}^{\pi}\cos^{2}\theta\sin\theta\,d\theta\,/\,\int_{0}^{\pi}\sin\theta\,d\theta=\frac{1}{3}$ represents the average projection of the dipole on the field direction for isotropic orientation in three dimensions, and $\left\langle|\vec{E}(\vec{r})|^{2}\right\rangle/|\vec{E}_{\mathrm{max}}|^{2}$ accounts for the average field intensity experienced by the emitters relative to the maximum. If the cavity mode field has the form $E(x)\sim\cos(kx)$, then the intensity is given by $|E(x)|^{2}\sim\cos^{2}(kx)=\left(1+\cos(2kx)\right)/2$. When averaging over the spatial coordinate $x$, we obtain $\langle\cos^{2}(kx)\rangle=1/2$. According to Eq. (14), this yields $\eta_{\mathrm{eff}}\sim 1/6$ as a reasonable estimate. Taking this factor into account, the realistic slope of the Lorentzian lifetime response curve becomes:
@@
[…]
From these considerations, it is clear that the proposed sensing mechanism can be considered as future-forward, targeting best-in-class performance, as it can potentially resolve $\Delta n$ changes down to $10^{-9}$ RIU. Even lower values can be readily achievable for cavities with $Q$-factors $>10^{7}$, achievable via state-of-the-art silicon photonics cavity designs [22]. This corresponds to a minimum detectable refractive index shift of $<10^{-9}$ RIU, outperforming plasmonic counterparts and unlocking single-molecule sensitivity in compact, integrated devices.

### v4

# Purcell-enhanced lifetime modulation of quantum emitters as a probe of local refractive index changes

@@ Abstract
Quantum emitters embedded in photonic integrated circuit (PIC) cavities offer a powerful and scalable platform for label-free refractive index sensing at the nanoscale. We propose and theoretically analyze a sensing mechanism based on Purcell-enhanced modulation of the emitter’s spontaneous emission lifetime, enabling detection of refractive index changes via time-correlated single-photon counting (TCSPC). In contrast to traditional resonance-shift sensors, our approach exploits the lifetime sensitivity to variations in the local density of optical states (LDOS), providing an intensity-independent, spectrally unresolvable, and potentially CMOS-compatible sensing modality. We derive analytical expressions linking refractive index perturbations to relative lifetime shifts and identify an optimal off-resonance operation regime where the lifetime response becomes linear and maximally sensitive to small perturbations. As a PIC material, silicon offers seamless integration of single-photon avalanche detectors suitable for TCSPC. We therefore illustrate the applicability of the proposed mechanism for quantum emitters in silicon. Parametric evaluation of the analytical expressions shows that, for moderate-quality photonic cavities ($Q=10^{5}$$10^{7}$), this method enables refractive index detection limits as low as $10^{-9}$ RIU — competitive with or even outperforming state-of-the-art plasmonic and microresonator sensors, yet requiring significantly simpler instrumentation. Furthermore, long-lived emitters such as T-centers in silicon offer a unique advantage, allowing sub-nanosecond lifetime shifts to be resolved with standard TCSPC systems. Although room-temperature operation of quantum emitters in silicon has yet to be demonstrated, our results lay the theoretical foundation for scalable, room-temperature, quantum-enabled refractive index sensing — eliminating the need for spectral resolution and cryogenic infrastructure. Given the generic nature of the proposed approach, and with ongoing advances in PIC technologies such as diamond-on-silicon and hybrid silicon/silicon nitride and silicon/silicon carbide platforms, the method can be readily extended to materials in which room-temperature operation of quantum emitters has already been experimentally demonstrated.
@@ I Introduction
However, their practical use at room temperature remains constrained by phonon-induced dephasing and non-radiative recombination, which broaden and suppress ZPL emission. To address these limitations, we propose a fundamentally different transduction approach: lifetime-based sensing via modulation of the local density of optical states (LDOS). Unlike conventional schemes based on spectral shifts, this method relies on variations in the emitter’s spontaneous emission rate induced by its electromagnetic environment. Such a detection scheme is inherently more robust to thermal broadening and inhomogeneous line shifts, and remains functional even when the ZPL becomes spectrally unresolved. In other words, even if the emitter rapidly loses coherence (i.e., short $T_{2}$) and exhibits a broad spectral line, the excited-state population may still decay on a much longer timescale ($T_{1}$), yielding a measurable lifetime — the quantity we exploit for sensing. Thermal effects can reduce the quantum efficiency of the emitter, and thus proportionally reduce the achievable sensitivity in our scheme; however, because the detection relies on spontaneous emission lifetime ($T_{1}$) rather than coherent optical dynamics ($T_{2}$), the method remains functional even in the presence of strong spectral broadening and dephasing. In this case, the sensitivity scales approximately with the room-temperature quantum efficiency of the emitter, meaning that a moderate reduction in quantum efficiency leads to a proportional, but not prohibitive, reduction in performance. This approach extends beyond T-centers to a broader class of T-like color centers, including group III — carbon complexes such as B–C, Al–C, and In–C [7, 8]. These centers are predicted to exhibit similar spin and optical properties, with higher structural symmetry (e.g., $C_{2v}$), improved radiative efficiencies, and greater thermal stability — making them promising candidates for scalable integration with silicon photonics.
@@
While our detection concept does not rely on spin coherence ($T_{2}$) and is based solely on the excited-state lifetime ($T_{1}$), demonstrations of exceptionally long $T_{2}$ in phosphorus donor nuclear spins — exceeding 39 minutes at room temperature in isotopically purified silicon [2] — and in 29Si nuclear spin qubits without isotopic enrichment [3] affirm silicon’s status as an ultra-low-noise host for quantum emitters. Analogous temperature studies of color centers such as SiV in diamond provide insights into emission linewidths: homogeneous broadening increases with temperature following a $T^{3}$ dependence, leading to ZPL broadening from $\sim$0.1–0.6 nm at cryogenic temperatures to $\sim$1.3–2 nm at room temperature [9]. A similar trend is anticipated for silicon-based centers, though experimental verification remains pending. Although room-temperature operation of quantum emitters in silicon remains unproven, no conclusive physical limitation has been established, suggesting potential for future advancement through photonic or material engineering.
@@
In light of these considerations, our proposed sensing scheme — based on photonic cavity-enhanced lifetime modulation of optically active color centers [10, 11, 12, 13, 14] — remains viable under room-temperature conditions provided that the emitter retains a measurable radiative decay channel. While thermal broadening will reduce the maximum attainable Purcell factor, the off-resonance operating regime and lifetime-based detection principle ensure that refractive index sensitivity scales linearly with the retained radiative efficiency. By decoupling sensing performance from spectral resolution constraints, it offers a robust and scalable platform for refractive index detection and single-molecule interaction studies. It is inherently compatible with the future discovery and optimization of silicon-based quantum emitters.
@@
However, it should be noted that, given the generic nature of the proposed approach and the ongoing advances in photonic integrated circuit (PIC) technologies — such as diamond-on-silicon, hybrid silicon/silicon nitride, and silicon/silicon carbide platforms — the method can be readily extended to materials in which room-temperature operation of quantum emitters has already been experimentally demonstrated, including diamond, silicon carbide [19], and silicon nitride [20, 21].
Figure 1 illustrates a possible multiplexed realization of the proposed lifetime-based sensing scheme.
Figure 1: Schematic illustration of a possible multiplexed PIC implementation of the proposed sensing approach, enabling simultaneous detection across multiple channels.
Light from a common source is split by an on-chip beam splitter into multiple waveguide channels, each coupled to a high-Q photonic cavity in the “interaction area”. These cavities host quantum emitters functionalized with different biorecognition layers (aptamers, antibodies, etc.), enabling parallel detection of distinct analytes. Analytes can be precisely delivered to the interaction areas by properly designed microfluidic channels. Binding events locally alter the refractive index, modulating the Purcell factor and hence the fluorescence lifetime of the emitters. The emitted photons are collected and directed to dedicated photodetectors for time-correlated single-photon counting, enabling simultaneous, label-free sensing across multiple channels on a single PIC platform. Because the emission originates from cavity-coupled quantum emitters, the platform is fully compatible with efficient photon extraction and does not suffer from the limitations typically associated with high-Q passive cavities.
@@ II Theoretical framework and results
$\tau=\frac{\tau_{int}}{F_{P}},$ (3)
where $\tau_{int}$ is the intrinsic spontaneous emission lifetime of the emitter in free space or in an unstructured (homogeneous) dielectric. We assume that the optical cavity used to modify the spontaneous emission lifetime of the emitter is optimized to achieve the maximum LDOS and Purcell enhancement. Local changes in the refractive index induce a spectral detuning of the cavity resonance. This detuning reduces the LDOS at the emitter’s location, thereby suppressing the spontaneous emission rate and resulting in an increase in the emitter’s lifetime. This lifetime shift can be detected using time-correlated single-photon counting (TCSPC), forming the basis for a label-free sensing modality. Given this, we estimate the effect of a refractive index change on the spontaneous emission lifetime of a quantum emitter coupled to a photonic cavity. A local change in the refractive index alters the effective index of the cavity mode, $n_{\mathrm{eff}}$, leading to a shift in cavity’s resonance angular frequency $\omega_{\mathrm{cav}}$. In the regime of small perturbations, the shift in the resonance frequency can be expressed using first-order perturbation theory as:
$\frac{\Delta\omega}{\omega_{\mathrm{cav}}}\approx-\frac{1}{2}\frac{\int\Delta\varepsilon(\vec{r})\left|\vec{E}(\vec{r})\right|^{2}dV}{\int\varepsilon(\vec{r})\left|\vec{E}(\vec{r})\right|^{2}dV},$ (4)
where $\Delta\varepsilon(\vec{r})=2n(\vec{r})\Delta n(\vec{r})$ is the permittivity perturbation induced by the local refractive index change $\Delta n(\vec{r})$, and $\vec{E}(\vec{r})$ is the cavity mode electric field. It should be noted here that in this work, we focus on refractive index variations arising from molecular binding events, while other possible $\Delta n$ sources (e.g., temperature fluctuations) are considered parasitic and should be minimized in practical implementations. By assuming $\Delta\varepsilon=2n\Delta n\approx 2n_{\mathrm{eff}}\Delta n$, we can then simplify Eq. (4) to:
@@
where $\Delta\omega=\omega_{\mathrm{emitter}}-\omega_{\mathrm{cav}}$ is the detuning. If the cavity is on-resonance in the initial state: $\Delta\omega=0$, and $F_{P}=F_{P,\mathrm{max}}$, and lifetime is at its minimum ($\tau=\tau_{int}/F_{P}$). Then, a small change in $F_{P}$ leads to the following relative lifetime change of the emitter:
@@
Figure 2: Simulated lifetime change ($\Delta\tau/\tau$) versus refractive index change ($\Delta n$) (i.e., sensitivity): (a) linear scale for $\Delta\tau/\tau$ from 0 to 10% and logarithmic scale for $\Delta n$; (b) logarithmic scale for $\Delta\tau/\tau$ from $10^{-1}$ to $10^{5}$ and logarithmic scale for $\Delta n$.
@@
However, at resonance ($\Delta\omega=0$), the Purcell factor is at its maximum, but its derivative with respect to detuning is zero. Therefore, in the immediate vicinity of $\Delta n=0$ (i.e., near the point $n_{\mathrm{eff}}$, see Figures 2a and 3), the dependence of $\Delta\tau/\tau$ on $\Delta n$ becomes non-linear (specifically, quadratic; see Eq. (7)), making it unsuitable for applications that require quantitative detection of small refractive index changes.
Figure 3: Simulated normalized Purcell factor versus refractive index change $\Delta n$: (a) linear scale for $\Delta n$ from $10^{-4}$ to $10^{-4}$; (b) log scale for $\Delta n$ from $10^{-9}$ to $10^{-3}$.
@@
We next derive the expression for the detection limit $\Delta n_{\mathrm{opt}}$ related to these optimal detection points for each $Q$ value. The emission lifetime $\tau$ at the shifted point $n_{0}$ can be written as:
@@
$=\pm 2\tau_{0}\frac{Q}{n_{\mathrm{eff}}}=\pm 4\tau_{\mathrm{eff}}\frac{Q}{n_{\mathrm{eff}}},$ (9)
here $\tau_{0}=\tau(n_{0})=2\tau(n_{\mathrm{eff}})=2\tau_{\mathrm{eff}}$. In the vicinity of $n_{0}$, we can write:
@@
Eq. (13) estimates the minimum detectable $\Delta n_{\text{opt}}$ for a given temporal resolution. According to Eq. (13), the detection limit of the proposed refractive index sensing approach in the off-resonance regime, $\Delta n_{\text{opt}}$, is limited by the $Q$ value of the photonic cavity and the resolution capability $\Delta\tau/\tau_{0}$ of the TCSPC system used. For instance, if a TCSPC system can resolve a 3% change in lifetime ($\Delta\tau/\tau_{0}=0.03$), then the minimum detectable $\Delta n_{\text{opt}}$ would be approximately $3.75\times 10^{-7}$ RIU for $Q=10^{5}$, $3.75\times 10^{-8}$ RIU for $Q=10^{6}$, and $3.75\times 10^{-9}$ RIU for $Q=10^{7}$, assuming $n_{\text{eff}}=2.5$. In addition, we calculated the linear dynamic range (LDR) for each $Q$ value as the span of refractive index changes $\Delta n$ over which the relative lifetime change $\Delta\tau/\tau_{0}$ remains approximately linear (with a deviation less than $\pm$10% from linear behavior) around the $n_{0}$ point. The LDR calculation details can be found in the Annex (see Supporting Information). The Python script and accompanying description for LDR calculation are available for download (see the Data Availability section at the end of the paper). These data are summarized in Table 2.
@@
$10^{5}$ $3.75\times 10^{-7}$ $1.12\times 10^{-5}$
$10^{6}$ $3.75\times 10^{-8}$ $1.12\times 10^{-6}$
$10^{7}$ $3.75\times 10^{-9}$ $1.12\times 10^{-7}$
@@
where $\left\langle\cos^{2}\theta\right\rangle=\int_{0}^{\pi}\cos^{2}\theta\sin\theta\,d\theta\,/\,\int_{0}^{\pi}\sin\theta\,d\theta=\frac{1}{3}$ represents the average projection of the dipole on the field direction for isotropic orientation in three dimensions, and $\left\langle|\vec{E}(\vec{r})|^{2}\right\rangle/|\vec{E}_{\mathrm{max}}|^{2}$ accounts for the average field intensity experienced by the emitters relative to the maximum. If the cavity mode field has the form $E(x)\sim\cos(kx)$, then the intensity is given by $|E(x)|^{2}\sim\cos^{2}(kx)=\left(1+\cos(2kx)\right)/2$ (here $k=\dfrac{2\pi}{\lambda}$ is the wavevector of the cavity mode and $\lambda$ is the wavelength of the mode). When averaging over the spatial coordinate $x$, we obtain $\langle\cos^{2}(kx)\rangle=1/2$. According to Eq. (14), this yields $\eta_{\mathrm{eff}}\sim 1/6$ as a reasonable estimate. Taking this factor into account, the realistic slope of the Lorentzian lifetime response curve becomes:
@@
$\eta_{\text{eff}}=\frac{1}{N}\int_{V}\rho(\vec{r})\left\langle\cos^{2}\theta\right\rangle\cdot\frac{|\vec{E}(\vec{r})|^{2}}{|\vec{E}_{\text{max}}|^{2}}\,dV,$ (17)
where $N=\int_{V}\rho(\vec{r})\,dV$ is the total number of emitters. This continuous formalism remains normalized between 0 and 1, and allows for accurate modeling of spatially varying emitter densities and orientation distributions, enabling more precise predictions in engineered photonic environments.
From a practical point of view, however — e.g., for measuring fluorescence emission — the following aspects are more important. The fluorescence intensity decays exponentially, and the decay rate differs for different emitters depending on their position and orientation. Therefore, it is essential to determine the fluorescence decay averaged over both the orientation and position of an ensemble of the emitters. The normalized fluorescence decay function $f(t)$, accounting for orientation and spatial averaging, can be expressed as:
$f(t)=\frac{\displaystyle\int_{0}^{\pi}\sin\theta\,d\theta\int_{0}^{\lambda/2}\cos^{2}(kx)\,\exp\left(-\frac{t}{\tau(\theta,x)}\right)dx}{\displaystyle\int_{0}^{\pi}\sin\theta\,d\theta\int_{0}^{\lambda/2}dx},$ (18)
where the position- and orientation-dependent lifetime $\tau(\theta,x)$ is defined as:
$\tau(\theta,x)=\frac{\tau_{\text{int}}}{F_{P}^{\text{max}}\cos^{2}(kx)\cos^{2}(\theta)}.$ (19)
Eq. (18) models the normalized fluorescence decay curve $f(t)$ for an ensemble of quantum emitters embedded in a cavity. Each emitter may be located at a different position within the cavity and oriented at a different angle with respect to the local electric field. These factors modulate the spontaneous emission rate due to the Purcell effect. The model accounts for this by averaging the decay function over both spatial and orientational distributions. Eq. (18) was solved numerically, and Figure 4 shows dependence of the normalized fluorescence decay $f(t)$ on the time $t/\tau_{int}$.
@@
According to Figure 4, the deviation of the averaged fluorescence decay curve from a straight line in logarithmic scale confirms its non-exponential nature. This behavior reflects the distribution of local lifetimes across the ensemble due to variations in emitter position and dipole orientation, resulting in a multi-exponential decay profile. Such effects are well known in complex photonic and biological systems [25], where spatial and structural inhomogeneities introduce deviation from a single-exponential decay and lead to a broadening of the decay dynamics.
While this may seem trivial — since even the sum of just two exponentials already deviates from a single-exponential function — it has practical implications. In particular, extracting lifetime information or classifying decay behavior becomes nontrivial when the signal contains a continuum of lifetime contributions. In this context, machine learning and AI-based tools are promising. They can be trained to identify patterns in TCSPC data, rapidly approximate multi-lifetime combinations, or directly classify decay signals without rigid model assumptions.
Furthermore, two practical aspects are important when interpreting experimental decay curves. First, while fluorescence signals are measured in absolute units (e.g., photon counts per time bin), the relative uncertainty increases as the signal decreases. This means that in the ”tail” of the decay, even small fluctuations in absolute counts can result in significant relative error. Second, as the decay slows at longer times — clearly visible in the logarithmic scale — larger time bin widths may be used (i.e., coarser binning), which helps reduce data volume. However, achieving sufficient signal-to-noise ratio in this region still requires extended acquisition time or higher photon collection efficiency.
To summarize, model provided by the Eq. (18) realistically predicts a non-exponential decay curve due to variations in emitter positions and orientations. It provides a physically accurate description of ensemble fluorescence in cavity-enhanced settings and is directly applicable to the interpretation of TCSPC experiments. It can also support fitting algorithms and machine learning-based lifetime analysis in nanophotonic biosensors.
From these considerations, it is clear that the proposed sensing mechanism can be considered as future-forward, targeting best-in-class performance, as it can potentially resolve $\Delta n$ changes down to $10^{-9}$ RIU. Even lower values can be readily achievable for cavities with $Q$-factors $10^{7}$, achievable via state-of-the-art silicon photonics cavity designs [26]. This corresponds to a minimum detectable refractive index shift of $10^{-9}$ RIU, outperforming plasmonic counterparts and unlocking single-molecule sensitivity in compact, integrated devices.